Gamma Industries has net income of $100,000, and it has 940,000 shares of common stock outstanding. The company's stock currently trades at $30 a share. Gamma is considering a plan in which it will use available cash to repurchase 15% of its shares in the open market at the current $30 stock price. The repurchase is expected to have no effect on net income or the company's P/E ratio. What will be its stock price following the stock repurchase? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Current EPS = Net Income/Number of shares

= 100,000/940,000

= $0.1063829 per share

P/E Ratio = Price per share/Earnings per share

= 30/0.1063829

= 282 times

EPS After repurchase = 100,000/799,000

= $0.12515644

Price after Repurchase = 0.12515644*282

= $35.2941

i.e. $35.29